YuzuEllie:

The equation 24 x 2 + 25 x − 47 a x − 2 = − 8 x − 3 − 53 a x − 2 is true for all values of x ≠ 2 a , where a is a constant. What is the value of a ? A) -16 B) -3 C) 3 D) 16

tax:

So in order to find the value of a, we can start by simplifying the equation: \[\frac{ 24x^2 + 25x - 47 }{ ax - 2 } = 7-8x - 3 - \frac{ 53 }{ ax -2 }\] Now, we'll eliminate the fractions: $$24x^2 + 25x - 47 = (-8x - 3)(ax - 2) - 53$$ Then we'll expand the right side of the equation by using foil and this will give youl: $$24x^2 + 25x - 47 = -8ax^2 + 16x - 3ax + 6 - 53$$ Then we'll combine the like terms: $$24x^2 + 25x - 47 = -8ax^2 + (16 - 3a)x - 47$$ FInally we'll equate the coefficients: $$24 = -8a$$ $$a = \frac{24}{-8}$$ $$a = -3$$ FInally we'll verify with the x coefficients: $$25 = 16 - 3a$$ $$25 = 16 - 3(-3)$$$$25 = 16 + 9$$ $$25 = 25$$ So your final answer is -3 (B)

YuzuEllie:

the anwser is A good job

tax:

You're joking.

YuzuEllie:

no dead Salsa

tax:

At least I tried :sob: